# METDrive: Multi-modal End-to-end Autonomous Driving with Temporal Guidance

### v2

@@ Abstract
Multi-modal end-to-end autonomous driving has shown promising advancements in recent work. By embedding more modalities into end-to-end networks, the system’s understanding of both static and dynamic aspects of the driving environment is enhanced, thereby improving the safety of autonomous driving. In this paper, we introduce METDrive, an end-to-end system that leverages temporal guidance from the embedded time series features of ego states, including rotation angles, steering, throttle signals, and waypoint vectors. The geometric features derived from perception sensor data and the time series features of ego state data jointly guide the waypoint prediction with the proposed temporal guidance loss function. We evaluated METDrive on the CARLA leaderboard benchmarks, achieving a driving score of $70\%$, a route completion score of $94\%$, and an infraction score of $0.78$.
@@ I Introduction
Multi-modal end-to-end autonomous driving systems have shown significant promise in enhancing the robustness and reliability of autonomous vehicles [1], [2]. Sensor fusion, which integrates data from multiple sources such as cameras and LiDAR, has become a cornerstone in advancing these systems [3], [4]. However, the raw data acquired from perception sensors frequently include a multitude of irrelevant objects that do not significantly contribute to the ego vehicle’s motion planning. Consequently, the output of an end-to-end model that relies on the fused features of these sensors may be adversely affected by the presence of such irrelevant detection[5], [6]. We show that existing end-to-end methodologies, which do not incorporate ego-related states, exhibit shortcomings and failure cases in certain scenarios. In this paper, we propose that the inclusion of additional ego-related features as guidance for the encoders can serve as a viable solution to mitigate this issue. Thus, we present METDrive, a novel end-to-end system that incorporates temporal guidance based on inputs from ego states.
@@
Based on the above insights, in this paper, our contributions are concluded as follows:
@@ II Related Work
In the field of autonomous driving, particularly in end-to-end learning systems, significant strides have been made in leveraging deep learning architectures to improve vehicle navigation and decision-making processes [8], [9], [10]. The following section reviews recent advances in model architectures for self-driving cars, focusing on their contributions and relevance to their proposed models.
@@ II-A Learning from All Vehicles
II-B Trajectory Prediction and Multi-Modal Fusion
The trajectory prediction has been a central focus in recent models, where the objective is to generate accurate waypoints for the future path of the vehicle. The TransFuser model, introduced by Chitta et al. [12], employs a multi-modal fusion transformer that combines RGB images and LiDAR data. This architecture allows the model to leverage the complementary nature of different sensor inputs, thus improving the accuracy of waypoint predictions. The model’s use of self-attention mechanisms within the transformer architecture is particularly effective in incorporating global context into the decision-making process.
@@ III Temporal Guidance for End-to-end Autonomous Driving
Perception Branch. The perception branch includes image and LiDAR encoders, both based on ResNet, which extract geometric features from sensor data [19]. An attention-based feature fusion block is used to integrate these features effectively [20]. Through the use of attention-based fusion, geometric features are processed in conjunction with temporal input from both image and LiDAR data, thereby ensuring consistency for subsequent fusion with time series data [21], [22].
@@
Similarly, the embedded vertical tokens $T_{y}\in\mathbb{R}^{1\times 3B\times B}$ is computed with $\sin{\theta},u$ and $\delta y$.
@@
$\begin{split}\mathcal{L}_{\mathrm{Temporal}}=\alpha(\left\|\mathbf{y}(\mathbf{[F_{g},F_{t}]}_{t=0}^{t=\frac{n}{2}})-\mathbf{\hat{y}}(\mathbf{[F_{g},F_{t}]}_{[0,...,\frac{n}{2}]})\right\|^{2}_{2})\\ +\beta(\left\|\mathbf{y}(\mathbf{[F_{g},F_{t}]}_{t=\frac{n}{2}}^{t=n}-\mathbf{\hat{y}}(\mathbf{[F_{g},F_{t}]}_{[\frac{n}{2},...,n]})\right\|^{2}_{2}),\end{split}$ (5)
@@ IV Experimental Results
In TransFuser++ [13], the training receipt contains a two-stage training, where the perception branch is trained with the corresponding losses first. Then a fine-tuning is performed on the checkpoint with all the losses. In our paper, we trained METDrive in one stage with both the perception branch and the temporal branch. With 320K training samples, we trained our model on a single Nvidia RTX 4090 24G at batch size 16.
Experimental Results. To evaluate the performance of our model, the CARLA Longest6 benchmarks are used during our evaluation [7]. Table I shows the performance comparison of the models in the evaluation, where our proposed system achieves a better driving score, route completion, and infraction score compared to the reproduced results from recent approaches. In Table I, RC means the route completion, which denotes the percentage of the route distance completed by an agent. IS means the infraction score that aggregates all tracked infractions triggered by an agent as a geometric series. DS means the driving score which is the product between the route completion and the infraction penalty.
@@ V Conclusion and Future Work
In this paper, we identify a common shortcoming in existing end-to-end systems for the CARLA simulator: the absence of guidance from additional modalities, which may lead to failure cases in long-term tasks. To address this issue, we propose to utilize ego-related temporal cues, such as rotation angles, steering, throttle signals, and waypoint vectors, to guide the geometric features derived from perception sensors in predicting waypoints. Consequently, we design a temporal guidance loss to monitor this integration process. To evaluate our system, we conducted experiments on the CARLA Longest6 benchmark, achieving a $70\%$ driving score, a $94\%$ route completion score and a $0.78$ infraction score.
@@
In light of these limitations, our future research routine will focus on the acquisition of a higher fidelity dataset. This dataset will be sourced from expert human drivers who navigate a diverse range of traffic scenarios. By incorporating data from skilled human operators, we anticipate a reduction in the aforementioned noise and a resultant enhancement in the quality of the input data. Additionally, we intend to develop a more robust framework that encompasses a broader spectrum of time series encoders and feature fusion tailored for ego state inputs. This enhanced approach is expected to mitigate current issues and improve the overall performance and reliability of the system under development.

### v3

# METDrive: Multimodal End-to-end Autonomous Driving with Temporal Guidance

@@ Abstract
Multimodal end-to-end autonomous driving has shown promising advancements in recent work. By embedding more modalities into end-to-end networks, the system’s understanding of both static and dynamic aspects of the driving environment is enhanced, thereby improving the safety of autonomous driving. In this paper, we introduce METDrive, an end-to-end system that leverages temporal guidance from the embedded time series features of ego states, including rotation angles, steering, throttle signals, and waypoint vectors. The geometric features derived from the perception sensor data and the time series features of ego state data jointly guide the waypoint prediction with the proposed temporal guidance loss function. We evaluated METDrive on the CARLA leaderboard benchmarks, achieving a driving score of $70\%$, a route completion score of $94\%$, and an infraction score of $0.78$.
@@ I Introduction
Multimodal end-to-end autonomous driving systems have shown significant promise in enhancing the robustness and reliability of autonomous vehicles [1], [2]. Sensor fusion, which integrates data from multiple sources such as cameras and LiDAR, has become a cornerstone in advancing these systems [3], [4]. However, the raw data acquired from the perception sensors frequently include a multitude of irrelevant objects that do not significantly contribute to the ego vehicle’s motion planning. Consequently, the output of an end-to-end model that relies on the fused features of these sensors may be adversely affected by the presence of such irrelevant detection[5], [6]. We show that existing end-to-end methodologies, which do not incorporate ego-related states, exhibit shortcomings and failure cases in certain scenarios. In this paper, we propose that the inclusion of additional ego-related features as guidance for the encoders can serve as a viable solution to mitigate this issue. Thus, we present METDrive, a novel end-to-end system that incorporates temporal guidance based on inputs from ego states.
@@
Based on the above insights, our contributions are concluded as follows:
@@ II Related Work
In the field of autonomous driving, particularly in end-to-end learning systems, significant strides have been made in leveraging deep learning architectures to improve vehicle navigation and decision-making processes [8], [9], [10], [11]. The following section reviews recent advances in model architectures for self-driving cars, focusing on their contributions and relevance to their proposed models.
@@ II-A Learning from All Vehicles
II-B Trajectory Prediction and Multimodal Fusion
The trajectory prediction has been a central focus in recent models, where the objective is to generate accurate waypoints for the future path of the vehicle. The TransFuser model, introduced by Chitta et al. [13], employs a multimodal fusion transformer that combines RGB images and LiDAR data. This architecture allows the model to leverage the complementary nature of different sensor inputs, thus improving the accuracy of waypoint predictions. The model’s use of self-attention mechanisms within the transformer architecture is particularly effective in incorporating global context into the decision-making process.
@@ III Temporal Guidance for End-to-end Autonomous Driving
Perception Branch. The perception branch includes image and LiDAR encoders, both based on ResNet, which extract geometric features from sensor data [20]. An attention-based feature fusion block is used to integrate these features effectively [21]. Through the use of attention-based fusion, geometric features are processed in conjunction with temporal input from both image and LiDAR data, thereby ensuring consistency for subsequent fusion with time series data [22].
@@
Similarly, the embedded vertical tokens $T_{y}\in\mathbb{R}^{1\times 3B\times B}$ is computed with $\sin{\theta},u$, and $\delta y$.
@@
$\begin{split}\mathcal{L}_{\mathrm{Temporal}}=\alpha(\left\|\mathbf{y}(\mathbf{[F_{g},F_{t}]}_{t=0}^{t=\frac{n}{2}})-\mathbf{\hat{y}}(\mathbf{[F_{g},F_{t}]})_{[0,...,\frac{n}{2}]}\right\|^{2}_{2})\\ +\beta(\left\|\mathbf{y}(\mathbf{[F_{g},F_{t}]}_{t=\frac{n}{2}}^{t=n})-\mathbf{\hat{y}}(\mathbf{[F_{g},F_{t}]})_{[\frac{n}{2},...,n]}\right\|^{2}_{2}),\end{split}$ (5)
@@ IV Experimental Results
In [14], the training receipt contains a two-stage training, where the perception branch is trained with the corresponding losses first. Then a fine-tuning is performed on the checkpoint with all the losses. We trained METDrive in one stage with both the perception branch and the temporal branch. With 320K training samples, we trained our model on a single Nvidia RTX 4090 24G at batch size 16.
@@
Experimental Results. To evaluate the performance of our model, the CARLA Longest6 benchmarks are used [7]. Table I shows the performance comparison of the models in the evaluation, where our proposed system achieves a better driving score, route completion, and infraction score compared to the reproduced results from the recent approaches. In Table I, RC means the Route Completion, which denotes the percentage of the route distance completed by an agent. IS means the Infraction Score that aggregates all tracked infractions triggered by an agent as a geometric series. DS means Driving Score, which is the product between the route completion and the infraction penalty.
@@ V Conclusion and Future Work
In this paper, we identify a common shortcoming in the existing end-to-end systems for the CARLA simulator: the absence of guidance from additional modalities, which may lead to failure cases in long-term tasks. To address this issue, we propose to leverage the ego-related temporal cues, such as rotation angles, steering, throttle signals, and waypoint vectors, to guide the geometric features derived from the perception sensors in predicting waypoints. Consequently, we design a temporal guidance loss to monitor this integration process. To evaluate our system, we conducted experiments on the CARLA Longest6 benchmark, achieving the driving score of $70\%$, route completion score of $94\%$, and infraction score of $0.78$.
@@
In light of these limitations, our future research routine will focus on the acquisition of a higher fidelity dataset. This dataset will be sourced from the expert human drivers who navigate a diverse range of traffic scenarios. By incorporating data from the skilled human operators, we anticipate a reduction in the aforementioned noise and a resultant enhancement in the quality of the input data. Additionally, we intend to develop a more robust framework that encompasses a broader spectrum of time series encoders and feature fusion tailored for ego state inputs. This enhanced approach is expected to mitigate current issues and improve the overall performance and reliability of the system under development.
Acknowledgements
Research reported in this publication was financially supported by RSF grant No. 24-41-02039.